∫0^∞e^x³ dx is equal to


A) 1/3Γ(1/3)


B) 1/2Γ(1/3)


C) Γ(1/3)           D) 3Γ(1/3)


1
Expert's answer
2021-08-23T04:51:34-0400
"\int_{0}^{\infty}e^{-x^3}dx=\frac{1}{3}\int_{0}^{\infty}e^{-y}y^{1/3-1}dy=\frac{1}{3}\Gamma\left(\frac{1}{3}\right)"

Answer: A